OpenStudy (anonymous):

You and your spouse decide that you want to have three children: two boys and a girl. What are your chances of achieving this if: (1) you don't care in what order the children appear? (2) if you want to have boy, boy, girl?

OpenStudy (anonymous):



OpenStudy (anonymous):

so, if you don't care its .5 * .5 *.5 If you do care, I think what you need to do is set it up as a Pr (A|B) question; where Pr (A|B) = Pr(A intersecting B)/Pr (B) where Pr (B) = Pr (2 boys and a girl), and A = BBG.